OpenStudy (anonymous):

A colony of bacteria grows at a rate proportional to its size. Initially, there are 500 bacteria present. The colony doubles in size every 3 hours. How many bacteria are present after 14 hours? I understand that the formulaa A=Ce^kt must be used but I am unsure about how to find each variable aand how it relates to finding the answer?

OpenStudy (anonymous):

\[500\times 3^{\frac{14}{3}}\] then a calculator

OpenStudy (anonymous):

you can do it your way if you like but it is much more cumbersome because first you have to find k and then you have to compute

OpenStudy (anonymous):

want to do it that way?

OpenStudy (anonymous):

Haha id much rather do it your way as it seems a lot simpler. However, our professor is requiring us to work it out painfully using the above method

OpenStudy (anonymous):

yes it is much much simpler and more accurate too

OpenStudy (anonymous):

ok C is easy, C = 500

OpenStudy (anonymous):

to find k note that in 3 hours it triples, so replace t by 3 and A by 1500 and solve \[1500=500e^{3k}\] \[3=e^{3k}\] \[3k=\ln(3)\] \[k=\frac{\ln(3)}{3}\]

OpenStudy (anonymous):

then you have your k, i guess you could use a calculator to find the decimal k = 0.3362

OpenStudy (anonymous):

now your formula is \[500e^{.3362t}\] and then replace t by 5 to get your answer

OpenStudy (anonymous):

awesome. Im trying to work it out and use your help as an answer key. Thank you very much

OpenStudy (anonymous):

yw